# Generalizable Image Repair for Robust Visual Autonomous Racing

### v1

@@ V-D Generative Image Repair
Generative models such as GANs [6], VAEs [7], and diffusion [8] have been widely adopted for image restoration. Models such as VAEs have reasonable repair capabilities but limited performance compared to more complex models like GANs and diffusion. CycleGAN [9] and pix2pix [10] support unpaired image translation, while denoising diffusion achieves state-of-the-art performance in high-quality synthesis [24]. For real-time image repair, efficient GAN variants such as ESRGAN [37] and FastGAN [38] optimize repair speed. Transformer models like SwinIR [39] enhance denoising through window-based self-attention. Neural filtering techniques [40] and self-supervised denoising [41] provide lightweight alternatives to image restoration. Unfortunately, diffusion models have significant computational complexity, with inference times that are unrealistic for real-time implementation, such as in our case study of autonomous racing. We observed that GANs strike a balance between significant repair capabilities and relatively low computational cost. This is especially the case in the presence of paired data and with the addition of our controller loss that optimizes for repaired images that maximize controller performance.
VI CONCLUSION AND DISCUSSION
This paper proposed an image repair framework for vision-based autonomous racing with enhanced robustness against visual corruptions from sensor noise, adverse weather, and lighting variations. By leveraging advanced generative adversarial models, our system restores corrupted images before control inference, effectively reducing distribution shift between training and testing environments [42, 43]. Extensive experiments demonstrate that our approach significantly improves control performance across diverse corruption types. In particular, CycleGAN with a controller-aligned loss function achieves the best overall stability and accuracy, confirming the benefits of integrating control-aware objectives into the repair process [44]. Our findings suggest that real-time image repair, when coupled with generative modeling and task-specific supervision, is a promising direction for improving the reliability of vision-based autonomous systems under challenging and unpredictable conditions.
@@ VI CONCLUSION AND DISCUSSION
Another important direction is to align the controller-based and human-based perspectives on image quality. For example, the high-frequency noise observed in Fig. 4 arises because the CNN controller has implicitly learned to rely on high-frequency cues, which can exacerbate overfitting to noise or artifacts [49, 50]. As a consequence, our repair method learns to reproduce these high-frequency features in the repaired images. Future work will mitigate this effect by encouraging the controller to focus on more robust, human-aligned features. For instance, integrating randomized smoothing [17] into the control pipeline can reduce sensitivity to high-frequency perturbations, thus promoting the use of lower-frequency and more visually pleasing features. By combining such controller designs with repair mechanisms that prioritize perceptual coherence, we hope to achieve a balance between control performance and visual fidelity in autonomous driving systems.
ACKNOWLEDGMENT
The authors gratefully acknowledge the support of the University of Florida AI Scholars Program.

### v2

@@ V-D Generative Image Repair
Generative models have been widely adopted for image restoration. Models such as VAEs have reasonable repair capabilities but limited performance compared to more complex models like GANs and diffusion. For real-time image repair, efficient GAN variants such as ESRGAN [37] and FastGAN [38] optimize repair speed. Transformer models like SwinIR [39] enhance denoising through window-based self-attention. Neural filtering techniques [40] and self-supervised denoising [41] provide lightweight alternatives to image restoration. Unfortunately, diffusion models have significant computational complexity, with inference times that are unrealistic for real-time implementation, such as in our case study of autonomous racing. We observed that GANs strike a balance between significant repair capabilities and relatively low computational cost. This is especially the case in the presence of paired data and with the addition of our controller loss that optimizes for repaired images that maximize controller performance.
VI DISCUSSION AND CONCLUSION
This paper proposed an image repair framework for vision-based autonomous control with enhanced robustness against visual corruptions from sensor noise, adverse weather, and lighting variations. By leveraging advanced generative adversarial models, our system restores corrupted images before control inference, effectively reducing distribution shift between training and testing environments [42, 43]. Extensive experiments demonstrate that our approach significantly improves control performance across diverse corruption types. In particular, CycleGAN with a controller-aligned loss function achieves the best overall stability and accuracy, confirming the benefits of integrating control-aware objectives into the repair process [44]. Our findings suggest that real-time image repair, when coupled with generative modeling and task-specific supervision, is a promising direction for improving the reliability of vision-based autonomous systems under challenging and unpredictable conditions.
@@ VI DISCUSSION AND CONCLUSION
Another important direction is to align the controller-based and human-based perspectives on image quality. For example, the high-frequency noise observed in Fig. 4 arises because the CNN controller has implicitly learned to rely on high-frequency cues, which can exacerbate overfitting to noise or artifacts [49, 50]. As a consequence, our repair method learns to reproduce these high-frequency features in the repaired images. Future work will mitigate this effect by encouraging the controller to focus on more robust, human-aligned features. For instance, integrating randomized smoothing [17] into the control pipeline can reduce sensitivity to high-frequency perturbations, thus promoting the use of lower-frequency and more visually pleasing features. By combining such controller designs with repair mechanisms that prioritize perceptual coherence, we hope to achieve a balance between control performance and visual fidelity in vision-based control systems.